# A Diffused Background from Axion-like Particles in the Microwave Sky

### v1

@@ 4.5 Sources of variation in the ALP Power Spectrum due to cluster astrophysics
The ALP background spectrum will depend on the single cluster power spectrums ($\alpha_{\ell}$’s), which depend on the astrophysics of the galaxy clusters via their electron density and magnetic field profiles. For all clusters considered, the profile parameters have been assigned random values within the allowed range. In principle, these would depend on the masses of clusters and their evolution at different redshifts.
Figure 7: (a) Variation of the ALP power spectrum with electron density strength $\mathrm{n_{e}}$. The power spectrum strength varies inversely with the square of electron density ($\mathrm{D_{\ell}^{\alpha\alpha}\propto\mathrm{n_{e}^{-2}}}$), when all mass ALPs are formed.(b) Variation of the ALP power spectrum with Magnetic field magnitude |B|. The power spectrum strength varies proportionally with the quadruple of electron density ($\mathrm{D_{\ell}^{\alpha\alpha}\propto\mathrm{\textbf{|B|}^{4}}}$). The CMB E-mode spectrum is also plotted for reference.
@@
Here we consider the variation of ALP diffused spectrum with a change in electron density and magnetic field strengths. With the mass range of $10^{-15}-10^{-11}$ eV, ALPs can be formed at every location for most of the galaxy clusters. Since the probability of conversion goes as $\mathrm{P(\gamma\rightarrow a)\propto\frac{1}{|\nabla n_{e}|}}$, the signal strength at every location within the cluster will vary as $\mathrm{\propto n_{e}^{-1}}$. Thus the background power spectrum shall vary as $\mathrm{D_{\ell}\propto{n_{e}^{-2}}}$ (see Fig. 7). Since the conversion probability $\mathrm{P(\gamma\rightarrow a)\propto|\textbf{B}|^{2}}$, with ALPs being formed at every location, the ALP background power spectrum will vary as $\mathrm{D_{\ell}\propto{|\textbf{B}|^{4}}}$ as shown in Fig.7. A higher magnetic field leads to a higher probability of conversion at all locations in the galaxy cluster.
@@ 5 Power spectrum estimation from sky maps
$\mathrm{\tilde{C_{\ell}^{i}}=B_{\ell}^{-2}\left[\frac{1}{2\ell+1}\sum_{m=-{\ell}}^{{\ell}}|a_{\ell m}^{obs}|^{2}-N_{\ell}\right]-\sum_{j\neq i}C_{\ell}^{j}.}$ (5.1)
@@
$\mathrm{Cov(\tilde{C_{\ell}})=\langle C_{\ell}^{2}\rangle-\langle C_{\ell}\rangle^{2}=\frac{2}{(2\ell+1)f_{sky}}[\sum_{i}C_{\ell}^{i}+B_{\ell}^{-2}N_{\ell}]^{2}.}$ (5.2)
@@
For a weighted combination of different frequency maps smoothed to common beam resolution, (see [31]) applied in various cleaning techniques such as ILC, the estimator is given as:
$\mathrm{\tilde{C_{\ell}^{i}}=B_{\ell}^{-2}\left[\frac{1}{2\ell+1}(\sum_{m=-l}^{l}\sum_{\nu}\sum_{\nu^{\prime}}w_{\nu}w_{\nu^{\prime}}a_{\ell m,\nu}^{obs\,*}a_{\ell m,\nu^{\prime}}^{obs})-\sum_{\nu}w_{\nu}^{2}N_{\ell}^{\nu}\right]-\sum_{j\neq i}\sum_{\nu}\sum_{\nu^{\prime}}w_{\nu}w_{\nu^{\prime}}\langle a_{\ell m,\nu}^{j\,*}a_{\ell m,\nu^{\prime}}^{j}\rangle.}$ (5.3)
Here we have considered the independence of various components.
@@ 6.1 ALP background as compared to galactic foregrounds
Synchrotron emission weakens with an increase in frequency, while dust increases with frequencies in the microwave and radio regions of the EM spectrum (Fig.8). ALP diffused spectrum increases with frequencies, following a $\mathrm{D_{\ell}\propto\nu^{2}I_{cmb}(\nu)^{2}}$ dependence, as compared to dust which follows a modified black-body spectrum. Both the galactic thermal dust and synchrotron emissions influence the power spectrum significantly at low multipoles, but at high multipoles, they weaken out, while the ALP diffused spectrum increases with multipoles following a $\mathrm{D_{\ell}\propto\ell^{2}}$ dependence. These spatial and spectral variations of the ALP signal as compared to CMB and foregrounds can be used to detect the diffused ALP background spectrum.
Figure 8: We show the spatial variation of the ALP background power spectrum ($\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$) at different multipoles ($\ell$) for various CMB-S4 frequency bands, as compared to CMB, galactic dust and galactic synchrotron emission. The CMB stays independent of frequency, while synchrotron weakens with an increase in frequency. Dust and ALP background spectrums increase with frequency, but they have a different multipole dependence, with dust power weakening at high multipoles, as opposed to ALP spectrum which increases at high multipoles. Also, dust follows the modified black-body model, while ALP spectrum simply scales as the squared of frequency.
@@ 6.2 SNR for different CMB surveys
The galaxy clusters are generated on the masked sky map in the unmasked regions. The sky fraction being observed depends on the detector, with sky fraction $\mathrm{f_{sky}=0.4}$ for SO, while 0.5 for CMB-S4 and CMB-HD. The ALP distortion signal is simulated in these clusters. Beam smearing (denoted by $\mathrm{B_{\ell}}$) occurs due to the resolution of the instrument and depends on the point source function. The combined map is then smeared with a Gaussian beam and instrument noise is added. The instrument noise for upcoming CMB surveys are taken assuming a Gaussian distribution.
We check the detectability of the diffused spectrum using current and future detectors. The CAMB [27] is used to generate the CMB power spectrum and map. The SNR is calculated taking into account the contributions from the optimum multipole range, using both the polarized maps for various frequencies of observation. Thus, the squared signal-to-noise ratio (SNR, denoted by $\rho$) for the distortion signal power spectrum is found by summing over the contributions from Q and U polarized maps for the multipole range $\mathrm{\ell_{min}}$ to $\mathrm{\ell_{max}}$. We use the values of $\mathrm{\ell_{min}}=1000$ and $\mathrm{\ell_{max}}$ corresponding to the beam resolution (multipole value at which $\mathrm{B_{\ell}^{2}=1/e}$) to obtain the SNR:
$\mathrm{\rho_{\nu}^{2}=\sum_{p=Q,U}\sum_{{\ell}_{min}}^{{\ell}_{max}}\frac{(C_{{\ell},1h}^{p,\nu}+C_{{\ell},2h}^{p,\nu})^{2}}{\frac{2}{(2{\ell}+1)f_{sky}}(C_{{\ell},cmb}^{p}+B_{{\ell},\nu}^{-2}N_{\ell}^{p,\nu})^{2}},}$ (6.1)
where $\mathrm{C_{\ell}}$ and $\mathrm{N_{\ell}}$ are the true signal and noise power spectrums respectively.
@@
Figure 9: This figure shows the SNR obtained at various frequency bands for different CMB detectors, corresponding to a coupling constant of $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$. The maxima for each detector is obtained for frequencies around 140 - 150 GHz, due to low beam size and instrumental noise. The presence of foregrounds reduces the SNR, but their impact is minute at frequencies 95 and 145 GHz, due to the weakening of the foregrounds in this frequency range.
Upcoming experiments such as the CMB-S4 and CMB-HD could be particularly useful in estimating this power spectrum with a lower beam and instrument noise. With regards to the current detectors, a coupling strength of $\mathrm{10^{-12}\,GeV^{-1}}$ with SO [4] configuration generates a maximum SNR of $\sim$ 0.52 (Fig. 9). The SNR achieves the highest value of 4.36 at the 145 GHz band in the presence of foregrounds for CMB-S4 [5], due to low noise and high resolution. A SNR of $\sim$ 93.87 will be achievable with CMB-HD [6] in the 150 GHz band. The presence of galactic foregrounds reduce the SNR, especially for low (impacted by synchrotron) and high frequency channels (impacted by dust). The bands from 90 to 170 GHz face the minimum impact as both dust and synchrotron weaken in this range of frequencies. The SNR varies as $\mathrm{g_{a\gamma}^{4}}$ with the coupling constant.
@@ 7 ILC to extract axion signal from multi-frequency maps
The ALP signal, contaminated with foregrounds and noise, accompanied by beam smearing, is difficult to extract. The spectral shape of the ALP background signal can be used to clean contaminants like CMB and foregrounds. The standard Internal Linear Combination (ILC) may be used to obtain a clean map as it minimizes the variance of the map and extracts a given spectrum. The higher the number of frequencies, accompanied with a lower beam size, the better the results obtained after ILC.
@@
$\mathrm{w_{ilc}=f_{a\gamma}^{T}C_{s}^{-1}(f_{a\gamma}^{T}C_{s}^{-1}f_{a\gamma})^{-1}.}$ (7.1)
Here $\mathrm{f_{a\gamma}}$ is the ALP distortion spectrum dependence on the frequency with $\mathrm{f_{a\gamma}\propto\nu I_{cmb}(\nu)}$. The ILC cleaned map is obtained as a weighted linear combination of the frequency maps ($\mathrm{S^{\nu}}$):
$\mathrm{S_{ilc}=\sum_{\nu}w_{ilc}^{\nu}S_{\nu}}.$ (7.2)
The distortion signal is simulated in galaxy clusters at different frequencies with a constant coupling constant ($\mathrm{g_{a\gamma}=\mathrm{1\times 10^{-12}\,GeV^{-1}}}$ here) for all clusters. The signal is then contaminated with CMB and galactic foregrounds. We have used PySM [32] to generate foreground maps. We use the synchrotron model "s-3" model which includes a curved index that flattens or steepens with frequency and dust model "d-3", which takes into account the spatial variation of spectral index on degree scales. The index is drawn from a Gaussian distribution. The various frequency maps are smoothed to a common beam resolution, given by the highest beam size among various frequency bands. The bands above 200 GHz can efficiently clean thermal dust contribution in the cleaned map. The CMB is frequency-independent, while the synchrotron emission is very weak above a frequency of 70 GHz. This lets us use just four bands with a higher beam resolution at frequencies greater than 90 GHz. Masking is done in galactic regions. The ALP signal dominates at high multipoles, owing to the one-halo term, while the foregrounds and CMB dampen out. It is the multipole range around $\ell=3000$ that can provide the best constraints on the ALP background spectrum, because at very high multipoles the noise takes over. Using ILC, we get the following SNRs for the various detectors for the ALP diffused spectrum with $\mathrm{z_{min}=1}$: SO - 0.24; CMB-S4 - 1.20; CMB-HD - 79.27. This SNR is the best we can achieve using ILC when the foregrounds are not well known. If they are well modelled, we can use template matching to achieve a higher SNR at the matched frequency bands corresponding to 140-150 GHz.
@@ 8 Constraining the ALP coupling constant using ILC
We perform the Bayesian estimation of the photon-ALP coupling constant from diffused ALP spectrum. The background spectrum scales as $\mathrm{g_{a\gamma}^{4}}$ (Fig.6). Since the electron density and magnetic field information for the unresolved clusters is lacking, and from Sec.4.5, we know that the background spectrum scales as $\mathrm{n_{e}^{-2}}$ and $\mathrm{|\textbf{B}|^{4}}$ (Fig.7), we thus obtain bounds on the quantity which we will call the modified coupling constant: $\mathrm{g_{a\gamma}^{2}|\textbf{B}|^{2}/n_{e}}$. The fiducial values we have used for the electron density and magnetic field strength parameters are: $\mathrm{n_{0}=10^{-3}\,cm^{-3}}$ and $\mathrm{B_{0}=0.1\,\mu G}$. We scale the modified coupling constant with respect to this choice of values to obtain bounds on the scaled quantity given as:
$\mathrm{A=\left(\frac{g_{a\gamma}}{10^{-12}\,GeV^{-1}}\right)^{2}\left(\frac{|\textbf{B}|}{0.1\,\mu G}\right)^{2}\left(\frac{n_{e}}{10^{-3}\,cm^{-3}}\right)^{-1}}.$ (8.1)
@@
Since the ALP background spectrum depends significantly on the minimum redshift $\mathrm{z_{min}}$ and minimum cluster mass $\mathrm{M_{min}}$, we obtain constraints on the modified constant for $\mathrm{z_{min}}=0.5$ and $\mathrm{z_{min}}=1$. For a given $\mathrm{z_{min}}$, we show the bounds for different choices of $\mathrm{M_{min}}$. We vary A with the upper bound set by the values: $\mathrm{g_{a\gamma}=10^{-11}\,GeV^{-1}}$, $\mathrm{n_{0}=0.5\times 10^{-3}\,cm^{-3}}$ and $\mathrm{B_{0}=0.5\,\mu G}$. We combine the posteriors from different mass ranges to obtain the bounds for a given choice of $\mathrm{z_{min}}$.
@@
We compare the bounds on modified coupling constant obtained using ILC for different CMB surveys: SO [4], CMB-S4 [5] and CMB-HD [6]. We simulate the fiducial maps at various frequencies without ALP signal ($\mathrm{g_{a\gamma}=0}$). We obtain the weights for these maps using Eq.7.1 and linearly combine them with their respective weights to obtain the ILC map. This method extracts the ALP signal, while minimizing the variance of the ILC map. We obtain the power spectrum of this map which gives us the term $\mathrm{\sum_{m=-\ell}^{\ell}|a_{lm}^{obs}|^{2}}$ in Eq.5.1 at different multipoles. Since noise for different frequency maps is not correlated, we obtain the term $\mathrm{N_{\ell}}$ in Eq.5.1 by combining weighted noise maps at various frequencies. Also, we obtain the mean power spectrum of different realizations of the combined fiducial ILC map (without ALP signal) by combining the fiducial maps at different frequencies (based on their ILC weights) to obtain the term $\mathrm{C_{\ell}^{cmb}+C_{\ell}^{fg}}$. These give us an estimation of the ALP background spectrum $\mathrm{C_{\ell}^{\alpha\alpha}}$. The beam $\mathrm{B_{\ell}}$ corresponds to the maximum beam size among the various frequency bands, as all maps are smoothed with the highest beam size before ILC weights are obtained.
Figure 10: Constraints from CMB-S4 using ILC for $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
@@
We use 50% partial sky to obtain bounds using CMB-S4 configuration, with frequency bands 95, 145, 220 and 270 GHz with the beam 2.2 arcmin corresponding to 95 GHz band. The bounds are stronger in case of a lower $\mathrm{M_{min}}$ as can be seen in Fig.10. In principle the improvement when using a lower $\mathrm{M_{min}}$ will depend on the relation between mass and electron density of low mass galaxy clusters, which will determine their contribution to the ALP background.
@@
The constraints are also tighter in the case of $\mathrm{z_{min}}=0.5$ as compared to $\mathrm{z_{min}}=1$ (Fig.10), as expected with low redshift clusters contributing to the background. Combining the posteriors for different mass ranges, we obtain the bounds on modified coupling constant using ILC with CMB-S4 configuration up to 95% confidence interval at $\mathrm{A<1.116}$ for $\mathrm{z_{min}=0.5}$, while $\mathrm{A<1.254}$ for $\mathrm{z_{min}=1}$.
Figure 11: Constraints from SO using ILC for $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
@@
For SO bounds, we use 40% sky fraction. The frequency bands 93, 145, 225, 280 GHz are used with the common beam 2.2 arcmin corresponding to the 93 GHz band. The constraints are shown in Fig.11. We obtain the following bounds using ILC, up to 95% confidence interval: $\mathrm{A<3.178}$ for $\mathrm{z_{min}=0.5}$; $\mathrm{A<3.724}$ for $\mathrm{z_{min}=1}$.
Figure 12: Converted ALP coupling constant constraints using ILC for various detectors with $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
@@
CMB-HD will have a partial sky fraction of 50%. We use the 90, 150, 220 and 280 GHz frequency bands to obtain bounds on the modified coupling constant with the common beam resolution of 0.42 arcmin, corresponding to the band 90 GHz. The results are shown in Fig.12. Using ILC, the bounds on modified coupling constant upto 95% confidence interval (C.I.) are: $\mathrm{A<0.115}$ for $\mathrm{z_{min}=0.5}$; $\mathrm{A<0.131}$ for $\mathrm{z_{min}=1}$. CMB-HD will provide bounds significantly better than CMB-S4 and SO on the modified coupling constant due to a higher beam resolution and improved sensitivity. The constraints obtained using the template matching of foregrounds are discussed in Sec.D.
@@
Figure 13: Converted bounds on ALP coupling constant for various surveys with $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$. Also the

### v3

@@ 4.5 Sources of variation in the ALP Power Spectrum due to cluster astrophysics
The ALP background spectrum will depend on the single cluster power spectra ($|\alpha_{\ell}|$’s), which depend on the astrophysics of the galaxy clusters via their electron density and magnetic field profiles. For all clusters considered, the profile parameters have been assigned random values within the allowed range. In principle, these would depend on the masses of clusters and their evolution at different redshifts.
Figure 7: (a) Variation of the ALP power spectrum with electron density strength $n_{e}$. The power spectrum strength varies inversely with the square of electron density ($D_{\ell}^{\alpha\alpha}\propto n_{e}^{-2}$), when all mass ALPs are formed.(b) Variation of the ALP power spectrum with magnetic field magnitude |B|. The power spectrum strength varies proportionally with the quadruple of magnetic field ($D_{\ell}^{\alpha\alpha}\propto\mathrm{\textbf{|B|}^{4}}$). The CMB E-mode spectrum is also plotted for reference.
@@
Here we consider the variation of ALP diffused spectrum with a change in electron density and magnetic field strengths. With the mass range of $10^{-15}-10^{-11}$ eV, ALPs can be formed at every location for most of the galaxy clusters. Since the probability of conversion goes as $P(\gamma\rightarrow a)\propto|\nabla n_{e}|^{-1}$, the signal strength at every location within the cluster will vary as $\propto n_{e}^{-1}$. Thus the background power spectrum shall vary as $D_{\ell}\propto{n_{e}^{-2}}$ (see Fig. 7). Since the conversion probability $P(\gamma\rightarrow a)\propto|\textbf{B}|^{2}$, with ALPs being formed at every location, the ALP background power spectrum will vary as $D_{\ell}\propto{|\textbf{B}|^{4}}$ as shown in Fig.7. A higher magnetic field leads to a higher probability of conversion at all locations in the galaxy cluster.
The polarized signal will be affected due to the scattering effects from regions of high-density electrons as the photon travels through the ICM of the clusters. Thus polarization information from high redshift clusters can be lost in the process. In our analysis, the primary contribution to the signal comes from low redshift (see Fig. 4 and Fig. 16 (in the appendix)). So, this effect will not impact the signal significantly. Another effect that can be important is the impact of the turbulent magnetic field and small-scale stochasticity on electron density. The presence of these effects will depolarise the signal and will produce an additional temperature fluctuation in the CMB due to ALPs [27]. In the future, we plan to study this effect from cluster simulations and by jointly predicting the ALP signature in both temperature and polarization anisotropy. The ellipticity and geometry of the clusters can also play a role, but since the background signal will be an integrated effect over many clusters, the impact due to variation in the shape will be averaged out. However, the shape of the galaxy cluster will be important for the search of ALP signatures from resolved galaxy clusters [34].
@@ 5 Power spectrum estimation from sky maps
$\tilde{C_{\ell}^{i}}=B_{\ell}^{-2}\left[\frac{1}{2\ell+1}\sum_{m=-{\ell}}^{{\ell}}|a_{\ell m}^{\mathrm{obs}}|^{2}-N_{\ell}\right]-\sum_{j\neq i}C_{\ell}^{j}.$ (5.1)
@@
$\mathrm{Cov}(\tilde{C_{\ell}})=\langle\tilde{C_{\ell}}^{2}\rangle-\langle C_{\ell}\rangle^{2}=\frac{2}{(2\ell+1)f_{\mathrm{sky}}}[\sum_{i}C_{\ell}^{i}+B_{\ell}^{-2}N_{\ell}]^{2}.$ (5.2)
@@ 6.1 ALP background as compared to galactic foregrounds
Synchrotron emission weakens with an increase in frequency, while dust increases with frequencies in the microwave and radio regions of the EM spectrum (Fig.8). ALP diffused spectrum increases with frequencies, following a $D_{\ell}\propto\nu^{2}I_{\mathrm{cmb}}(\nu)^{2}$ dependence, as compared to dust which follows a modified black-body spectrum. Both the galactic thermal dust and synchrotron emissions influence the power spectrum significantly at low multipoles, but at high multipoles, they weaken out, while the ALP diffused spectrum increases with multipoles following a $D_{\ell}\propto\ell^{2}$ dependence. These spatial and spectral variations of the ALP signal as compared to CMB and foregrounds can be used to detect the diffused ALP background spectrum.
Figure 8: We show the spatial variation of the ALP background power spectrum ($g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$) at different multipoles ($\ell$) for various CMB-S4 frequency bands, as compared to CMB, galactic dust and galactic synchrotron emission. The CMB stays independent of frequency, while galactic synchrotron emission weakens with an increase in frequency. Dust and ALP background spectra increase with frequency, but they have a different multipole dependence, with dust power weakening at high multipoles, as opposed to ALP spectrum which increases at high multipoles. Also, dust follows the modified black-body model, while ALP spectrum simply scales as the squared of frequency.
@@ 6.2 SNR for different CMB surveys
The galaxy clusters are generated on the masked sky map in the unmasked regions. The sky fraction being observed depends on the detector, with sky fraction $f_{\mathrm{sky}}=0.4$ for SO, while 0.5 for CMB-S4 and CMB-HD. The ALP distortion signal is simulated in these clusters. Beam smearing (denoted by $B_{\ell}$) occurs due to the resolution of the instrument and depends on the point source function. The combined map is then smeared with a Gaussian beam and instrument noise is added. The instrument noise for upcoming CMB surveys are taken assuming a Gaussian distribution.
We check the detectability of the diffused spectrum using current and future detectors. The CAMB [46] is used to generate the CMB power spectrum and map. The SNR is calculated taking into account the contributions from the optimum multipole range, using both the polarized maps for various frequencies of observation. Thus, the squared signal-to-noise ratio (SNR, denoted by $\rho$) for the distortion signal power spectrum is found by summing over the contributions from Q and U polarized maps for the multipole range $\ell_{\mathrm{min}}$ to $\ell_{\mathrm{max}}$. We use the values of $\ell_{\mathrm{min}}=1000$ and $\ell_{\mathrm{max}}$ corresponding to the beam resolution (multipole value at which $B_{\ell}^{2}=1/e$) to obtain the SNR:
$\rho_{\nu}^{2}=\sum_{p=\mathrm{Q,U}}\sum_{{\ell}_{\mathrm{min}}}^{{\ell}_{\mathrm{max}}}\frac{(C_{{\ell},1h}^{p,\nu}+C_{{\ell},2h}^{p,\nu})^{2}}{\frac{2}{(2{\ell}+1)f_{\mathrm{sky}}}(C_{{\ell},\mathrm{cmb}}^{p}+B_{{\ell},\nu}^{-2}N_{\ell}^{p,\nu})^{2}},$ (6.1)
where $C_{\ell}$ and $N_{\ell}$ are the true signal and noise power spectra respectively.
@@
Figure 9: This figure shows the SNR obtained at various frequency bands for different CMB detectors, corresponding to a coupling constant of $g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$. The maxima for each detector is obtained for frequencies around 140 - 150 GHz, due to low beam size and instrumental noise. The presence of foregrounds reduces the SNR, but their impact is minute at frequencies 95 and 145 GHz, due to the weakening of the foregrounds in this frequency range.
Upcoming experiments such as the CMB-S4 and CMB-HD could be particularly useful in estimating this power spectrum with a lower beam and instrument noise. With regards to the current detectors, a coupling strength of $10^{-12}\,\mathrm{GeV^{-1}}$ with SO [4] configuration generates a maximum SNR of $\sim$ 0.52 (Fig. 9). The SNR achieves the highest value of 4.36 at the 145 GHz band in the presence of foregrounds for CMB-S4 [5], due to low noise and high resolution. An SNR of $\sim$ 93.87 will be achievable with CMB-HD [6] in the 150 GHz band. The presence of galactic foregrounds reduces the SNR, especially for low (impacted by synchrotron) and high frequency channels (impacted by dust). The bands from 90 to 170 GHz face the minimum impact as both dust and synchrotron weaken in this range of frequencies. The SNR varies as $g_{a\gamma}^{4}$ with the coupling constant.
@@ 7 ILC to extract axion signal from multi-frequency maps
The ALP signal, contaminated with foregrounds and noise, accompanied by beam smearing, is difficult to extract. The spectral shape of the ALP background signal can be used to clean contaminants like CMB and foregrounds. The standard Internal Linear Combination (ILC) may be used to obtain a clean map as it minimizes the variance of the map and extracts a given spectrum. The higher the number of frequencies, accompanied by a lower beam size, the better the results obtained after ILC.
@@
$w_{\mathrm{ilc}}=f_{a\gamma}^{T}C_{s}^{-1}(f_{a\gamma}^{T}C_{s}^{-1}f_{a\gamma})^{-1}.$ (7.1)
Here $f_{a\gamma}$ is the ALP distortion spectrum dependence on the frequency with $f_{a\gamma}\propto\nu I_{\mathrm{cmb}}(\nu)$. The ILC cleaned map is obtained as a weighted linear combination of the frequency maps ($S^{\nu}$):
$S_{\mathrm{ilc}}=\sum_{\nu}w_{\mathrm{ilc}}^{\nu}S_{\nu}.$ (7.2)
The distortion signal is simulated in galaxy clusters at different frequencies with a constant coupling constant ($g_{a\gamma}=\mathrm{1\times 10^{-12}\,GeV^{-1}}$ here) for all clusters. The signal is then contaminated with CMB and galactic foregrounds. We have used PySM [55] to generate foreground maps. We use the synchrotron model "s-3" model which includes a curved index that flattens or steepens with frequency and dust model "d-3", which takes into account the spatial variation of spectral index on degree scales. The index is drawn from a Gaussian distribution. The various frequency maps are smoothed to a common beam resolution, given by the highest beam size among various frequency bands. The bands above 200 GHz can efficiently clean thermal dust contribution in the cleaned map. The CMB is frequency-independent, while the synchrotron emission is very weak above a frequency of 70 GHz. This lets us use just four bands with a higher beam resolution at frequencies greater than 90 GHz. Masking is done in galactic regions. The ALP signal dominates at high multipoles, owing to the one-halo term, while the foregrounds and CMB dampen out. It is the multipole range around $\ell=3000$ that can provide the best constraints on the ALP background spectrum, because at very high multipoles the noise takes over. Using ILC, we get the following SNRs for the various detectors for the ALP diffused spectrum with $z_{\mathrm{min}}=1$: 0.24 with SO; 1.20 with CMB-S4; 79.27 with CMB-HD. This SNR is the best we can achieve using ILC when the foregrounds are not well known. If they are well modelled, we can use template matching to achieve a higher SNR at the matched frequency bands corresponding to 140-150 GHz.
The ALP signal will be correlated with other weak signals like the CIB emission and polarized SZ effect. This correlation may affect the efficiency of the ILC cleaning method. To deal with these, we can also make use of methods like the constrained ILC (both for temperature and polarization fluctuations) [56, 49, 57], although at a noise tradeoff, which can remove contaminations with a given spectrum. It can make use of the spectral information of the contaminant to completely or partially eliminate it. We assume in our analysis that the effect of these contaminants is low and the spectral and spatial shape of the ALP spectrum can distinguish the signal from these contaminants.
@@ 8 Constraining the ALP coupling constant using ILC
We perform the Bayesian estimation of the photon-ALP coupling constant from diffused ALP spectrum. The background spectrum scales as $g_{a\gamma}^{4}$ (Fig.6). Since the electron density and magnetic field information for the unresolved clusters is lacking, and from Sec.4.5, we know that the background spectrum scales as $n_{e}^{-2}$ and $|\textbf{B}|^{4}$ (Fig.7), we thus obtain bounds on the quantity which we will call the modified coupling constant: $g_{a\gamma}^{2}|\textbf{B}|^{2}/n_{e}$. The fiducial values we have used for the electron density and magnetic field strength parameters are: $n_{0}=10^{-3}\,\mathrm{cm^{-3}}$ and $B_{0}=0.1\,\mathrm{\mu G}$. We scale the modified coupling constant with respect to this choice of values to obtain bounds on the scaled quantity given as:
$A=\left(\frac{g_{a\gamma}}{10^{-12}\,\mathrm{GeV^{-1}}}\right)^{2}\left(\frac{|\textbf{B}|}{0.1\,\mathrm{\mu G}}\right)^{2}\left(\frac{n_{e}}{10^{-3}\,\mathrm{cm^{-3}}}\right)^{-1}.$ (8.1)
@@
Since the ALP background spectrum depends significantly on the minimum redshift $z_{\mathrm{min}}$ and minimum cluster mass $M_{\mathrm{min}}$, we obtain constraints on the modified constant for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$. For a given $z_{\mathrm{min}}$, we show the bounds for different choices of $M_{\mathrm{min}}$. We vary $A$ with the upper bound set by the values: $g_{a\gamma}=10^{-11}\,\mathrm{GeV^{-1}}$, $n_{0}=0.5\times 10^{-3}\,\mathrm{cm^{-3}}$ and $B_{0}=0.5\,\mathrm{\mu G}$. We combine the posteriors from different mass ranges to obtain the bounds for a given choice of $z_{\mathrm{min}}$.
We compare the bounds on modified coupling constant obtained using ILC for different CMB surveys: SO [4], CMB-S4 [5] and CMB-HD [6]. Along with these experiments, dedicated CMB spectral distortion experiments such as the Primordial Inflation Explorer (PIXIE) [58, 59, 60] with high sensitivities and a high spectral resolution can also be used for measuring the coupling constant of ALPs using diffused background signal if the distortion polarization information is well resolvable. This will require a high beam resolution, otherwise the signal will be depolarized due to convolution with the beam. Future CMB experiments with an appropriate design with both spectral and spatial resolution can bring new discoveries in this area.
@@
We simulate the fiducial maps at various frequencies without ALP signal ($g_{a\gamma}=0$). We obtain the weights for these maps using Eq.7.1 and linearly combine them with their respective weights to obtain the ILC map. This method extracts the ALP signal, while minimizing the variance of the ILC map. We obtain the power spectrum of this map which gives us the term $\sum_{m=-\ell}^{\ell}|a_{lm}^{\mathrm{obs}}|^{2}$ in Eq.5.1 at different multipoles. Since noise for different frequency maps is not correlated, we obtain the term $N_{\ell}$ in Eq.5.1 by combining weighted noise maps at various frequencies. Also, we obtain the mean power spectrum of different realizations of the combined fiducial ILC map (without ALP signal) by combining the fiducial maps at different frequencies (based on their ILC weights) to obtain the term $C_{\ell}^{\mathrm{cmb}}+C_{\ell}^{\mathrm{fg}}$. These give us an estimation of the ALP background spectrum $C_{\ell}^{\mathrm{ax}}$. The beam $B_{\ell}$ corresponds to the maximum beam size among the various frequency bands, as all maps are smoothed with the highest beam size before ILC weights are obtained.
Figure 10: Constraints from CMB-S4 using ILC for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$
@@
We use 50% partial sky to obtain bounds using CMB-S4 configuration, with frequency bands 95, 145, 220 and 270 GHz with the beam 2.2 arcmin corresponding to 95 GHz band. The bounds are stronger in case of a lower $M_{\mathrm{min}}$ as can be seen in Fig.10. In principle the improvement when using a lower $M_{\mathrm{min}}$ will depend on the relation between mass and electron density of low mass galaxy clusters, which will determine their contribution to the ALP background.
@@
The constraints are also tighter in the case of $z_{\mathrm{min}}=0.5$ as compared to $z_{\mathrm{min}}=1$ (Fig.10), as expected with low redshift clusters contributing to the background. Combining the posteriors for different mass ranges, we obtain the bounds on modified coupling constant using ILC with CMB-S4 configuration up to 95% confidence interval at $A<1.116$ for $z_{\mathrm{min}}=0.5$, while $A<1.254$ for $z_{\mathrm{min}}=1$.
Figure 11: Constraints from SO using ILC for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$
@@
For SO bounds, we use 40% sky fraction. The frequency bands 93, 145, 225, 280 GHz are used with the common beam 2.2 arcmin corresponding to the 93 GHz band. The constraints are shown in Fig.11. We obtain the following bounds using ILC, up to 95% confidence interval: $A<3.178$ for $z_{\mathrm{min}}=0.5$; $A<3.724$ for $z_{\mathrm{min}}=1$.
Figure 12: Constraints from CMB-HD using ILC for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$.
@@
CMB-HD will have a partial sky fraction of 50%. We use the 90, 150, 220 and 280 GHz frequency bands to obtain bounds on the modified coupling constant with the common beam resolution of 0.42 arcmin, corresponding to the band 90 GHz. The results are shown in Fig.12. Using ILC, the bounds on modified coupling constant upto 95% confidence interval (C.I.) are: $A<0.115$ for $z_{\mathrm{min}}=0.5$; $A<0.131$ for $z_{\mathrm{min}}=1$. CMB-HD will provide bounds significantly better than CMB-S4 and SO on the modified coupling constant due to a higher beam resolution and improved sensitivity. The constraints obtained using the template matching of foregrounds are discussed in Sec.D.
@@
Figure 13: Converted bounds on ALP coupling constant for various surveys with $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$